What is the x-intercept for the following equation y=2(x+1)^2 + 3? Please show step by step.
Answer by ikleyn(52778)

.
What is the x-intercept for the following equation y=2(x+1)^2 + 3? Please show step by step.
~~~~~~~~~~~~~~~


            In this post,   y = 2(x+1)^2 + 3   IS  NOT   an equation.

            This equality is the  DEFINITION  of function  y.

            Use the terms correctly,  precisely with their destiny and meaning.


The quadratic term   2*(x+1)^2  is always non-negative.


With the constant term  "3"  added,  the function  y = 2*(x+1)^2 + 3  is always positive.


So the graph of the function  y = 2(x+1)^2 + 3  HAS NO  x-intercept.    ANSWER


Solved and explained.